# Lossy Beyond Diagonal Reconfigurable Intelligent Surfaces: Modeling and Optimization

### v1

@@ Abstract
Beyond diagonal reconfigurable intelligent surface (BD-RIS) has emerged as an advancement and generalization of the conventional diagonal RIS (D-RIS) by introducing tunable interconnections between RIS elements, enabling smarter wave manipulation and enlarged coverage. While BD-RIS has demonstrated advantages over D-RIS in various aspects, most existing works rely on the assumption of a lossless model, leaving practical considerations unaddressed. This paper thus proposes a lossy BD-RIS model and develops corresponding optimization algorithms for various BD-RIS-aided communication systems. First, by leveraging admittance parameter analysis, we model each tunable admittance component based on a lumped circuit with losses and derive an expression of a circle characterizing the real and imaginary parts of each tunable admittance. We then consider the received signal power maximization in single-user single-input single-output (SISO) systems with the proposed lossy BD-RIS model. To solve the formulated challenging optimization problem, we design an effective algorithm by carefully exploiting the problem structure. In particular, an alternating direction method of multipliers (ADMM) framework is custom-designed to deal with the complicated constraints associated with lossy BD-RIS. Furthermore, we extend the proposed algorithmic framework to more general multiuser multiple-input single-output (MU-MISO) systems, where the transmit precoder and BD-RIS scattering matrix are jointly designed to maximize the sum-rate of the system. Finally, simulation results demonstrate that all BD-RIS architectures still outperform D-RIS in the presence of losses, but the optimal BD-RIS architectures in the lossless case are not necessarily optimal in the lossy case, e.g. group-connected BD-RIS can outperform fully- and tree-connected BD-RISs in SISO systems with relatively high losses at BD-RIS, whereas the opposite always holds true in the lossless case.
Index Terms:
Beyond diagonal reconfigurable intelligent surface (BD-RIS), lossy RIS model, alternating direction method of multipliers.
@@ I Introduction
Reconfigurable intelligent surface (RIS) is a nearly-passive planar surface composed of numerous reflecting elements, each capable of inducing a phase and/or amplitude change to the impinging wave. This flexibility enables dynamic wave manipulation, which provides enhanced channel gain and enlarges wireless coverage [1, 2, 3]. According to microwave network theory [4], RIS can be generally modeled as numerous scattering elements connected to a multiport reconfigurable impedance network [5]. Most existing studies focus on RIS designs with a single-connected architecture, where each port of the reconfigurable impedance network is connected to the ground via a tunable impedance. This mathematically results in a diagonal scattering matrix [5], and such a single-connected RIS is also referred to as diagonal (D)-RIS. Recently, a novel branch of RIS, known as beyond diagonal (BD)-RIS [6], has been proposed as a generalization of D-RIS. BD-RIS operates by introducing additional tunable impedance components that interconnect the ports in the reconfigurable impedance network, mathematically resulting in a scattering matrix that is no longer restricted to being diagonal. This enhanced flexibility, provided by those interconnections, enables a smarter wave manipulation, leading to greater channel gain, higher spectral efficiency, and broader wireless coverage compared to D-RIS [6].
@@
Existing studies on BD-RIS have shown its advantages over D-RIS in various aspects. These works can be categorized as follows: [7, 8, 9] focus on mode analysis; [5, 10, 11, 12, 13] focus on architecture design; [14, 15, 16] focus on signal processing; [17, 18, 19, 20, 21] focus on hardware impairments. However, most existing works [5, 7, 8, 9, 10, 11, 14, 15, 13, 16, 17, 18, 19, 21, 12] assume the reconfigurable impedance network is lossless to simplify the theoretical analysis and explore the performance upperbound of BD-RIS architectures. One exception work [20] models BD-RIS with lossy interconnections in the reconfigurable impedance network using transmission line theory, where these interconnections are in series with tunable impedance components. However, the tunable impedance components themselves are still assumed to be lossless. In practice, semiconductor devices such as positive-intrinsic-negative (PIN) diodes and varactor diodes are used to implement the tunable impedance components. These semiconductor devices have inherent resistances, which dissipate power in the circuit. Therefore, assuming a reconfigurable impedance network with lossless tunable components results in a model that underestimates the performance degradation in realistic systems. It is crucial to analyze the characteristics of BD-RIS with lossy impedance components in the reconfigurable impedance network and establish an accurate model that reflects these losses.
@@
In the works on practical D-RIS modeling [22, 23, 24], a lumped circuit model for each RIS element has been shown to effectively capture the relationship between reflection amplitude and phase. However, generalizing this approach to lossy BD-RIS is challenging. D-RISs are typically modeled using scattering parameters, where each entry in the scattering matrix represents the amplitude and phase response of an individual RIS element, allowing them to be modeled independently. In contrast, for BD-RISs, the interconnections between RIS elements cause the amplitude and phase of each entry in the scattering matrix to depend on other entries as well, making it impossible to model individual entries independently. In [25], a universal framework has been proposed that unifies the analysis of scattering, impedance, and admittance parameters for characterizing RIS-aided models, and presents the advantages of each type of parameter analysis in RIS modeling and optimization. While the aforementioned works [5, 17, 14, 8, 9, 10, 11, 15, 13] primarily focus on modeling and designing BD-RIS-aided systems using scattering parameter, admittance parameter has proven to be a powerful approach in characterizing the circuit topology models of BD-RIS. Admittance parameter analysis enables the modeling of BD-RIS with lossy interconnections using transmission line theory [20], the wideband modeling of BD-RIS through circuit-based approaches [21], and the development of an efficient architecture-independent optimization framework for BD-RIS in general multiuser systems [16].
Moreover, from the optimization perspective, lossy BD-RIS introduces additional complex constraints compared to both practical D-RIS and lossless BD-RIS, due to the beyond diagonal and lossy properties, respectively. These complexities render existing D-RIS and BD-RIS optimization methods [5, 7, 17, 14, 8, 9, 11, 10, 15, 13, 25, 18, 19, 20, 21, 16, 22, 23, 24, 12] inapplicable.
In this paper, we address the research gap in modeling BD-RIS with lossy tunable impedance components. To tackle the above challenges, we leverage the advantages of admittance parameter analysis as in [20, 21, 16] to model and optimize BD-RIS with a reconfigurable admittance network that accounts for lossy tunable admittance components. The contributions of this work are summarized as follows.
@@
First, we propose a novel model for BD-RIS that incorporates losses in the reconfigurable admittance network. Specifically, by utilizing admittance parameter analysis, we represent the lossy reconfigurable admittance network by modeling each tunable admittance component individually using a lumped circuit model based on a practical varactor. In the proposed model, the real and imaginary parts of each admittance component form a circle, with feasible values constrained to a segment of the circle determined by the practical range of the tunable capacitance. The proposed model is sufficiently general to be applicable to any BD-RIS architectures, such as fully-/group-connected [5] and tree-/forest-connected architectures [^1] The proposed model can also be applied to band-connected and stem-connected architectures [12]. [10].
@@
Second, we optimize the scattering matrix of BD-RIS to maximize the received signal power for a BD-RIS-aided single-user single-input single-output (SISO) system, incorporating the proposed lossy BD-RIS model. To solve this problem, we employ a minorization-maximization (MM) framework to transform the problem into a more tractable form. To deal with the complicated constraints introduced by lossy BD-RIS, we further custom-design an alternating direction method of multipliers (ADMM) algorithm to solve the resulting subproblems. In addition, to reduce the computational complexity, we propose a low-complexity solution that minimizes the Euclidean distance of an admittance matrix condition derived from the lossless BD-RIS model, using a similar ADMM algorithm to handle the constraints introduced by lossy BD-RIS.
@@
Fourth, we present simulation results to evaluate the performance of the proposed algorithms. Results show that all BD-RIS architectures outperform D-RIS under the lossy RIS modeling. In contrast to the lossless case where fully- and tree-connected BD-RISs always outperform group-connected BD-RIS in SISO systems, in the lossy case, group-connected BD-RIS can outperform fully- and tree-connected BD-RISs. Except for the fully-connected architecture, group-connected BD-RIS consistently outperforms forest-connected BD-RIS under the same group size [^2] Recall that in lossless BD-RIS, forest-connected and group-connected architectures achieve the same performance for a given group size in SISO systems [10, 26], though the latter has higher complexity. In lossy systems, the greater number of interconnections in the group-connected architecture offers additional flexibility to combat losses, resulting in enhanced performance compared to the forest-connected architecture.. In MU-MISO systems, fully-connected BD-RIS maintains the best performance, which is consistent with the lossless case.
@@
Notations: Boldface lowercase and uppercase letters represent column vectors and matrices, respectively, while scalars are denoted with letters not in bold font. $\mathbb{R}$ and $\mathbb{C}$ indicate the set of real numbers and complex numbers, respectively. $(\cdot)^{*}$, $(\cdot)^{T}$, $(\cdot)^{H}$, and $(\cdot)^{-1}$ represent the conjugate, transpose, conjugate transpose, and inverse, respectively. ${}\Re\{\cdot\}$ and ${}\Im\{\cdot\}$ indicate the real and imaginary parts of a complex number, respectively. $\mathbb{E}\{\cdot\}$ denotes the statistical expectation. ||$\cdot$|| ⋅ |, $\|\cdot\|_{2}$ and $\|\cdot\|_{F}$ refer to the absolute-value norm, the $\ell_{2}$ norm, and the Frobenius norm, respectively. $\arg(\cdot)$ denotes the phase of a complex number. $\otimes$ represents the Kronecker product. $\mathsf{blkdiag}(\cdot)$ represents a block-diagonal matrix. $\mathsf{vec}(\cdot)$ denotes the vectorization of a matrix, $\overline{\mathsf{vec}}(\cdot)$ is the reverse operation of vectorization, and $\mathsf{tr}(\cdot)$ denotes the trace of a matrix. $\mathbf{I}_{M}$ denotes an $M\times M$ identity matrix. $\jmath=\sqrt{-1}$ represents the imaginary unit. $[\mathbf{A}]_{i,j}$ and $[\mathbf{a}]_{i}$ indicate the $(i,j)$-th entry of $\mathbf{A}$ and the $i$-th entry of $\mathbf{a}$, respectively. Finally, $[\mathbf{A}]_{i:i^{\prime},j:j^{\prime}}$ represents the submatrix of $\mathbf{A}$ formed by selecting the $i$-th to $i^{\prime}$-th rows and $j$-th to $j^{\prime}$-th columns.
@@ II-A Admittance Parameter Analysis of BD-RIS
An $M$-element RIS is a passive device, which is modeled as $M$ elements connected to an $M$-port reconfigurable admittance network [5]. The reconfigurable admittance network, composed of tunable passive components, can be mathematically described by its admittance matrix $\mathbf{Y}\in\mathbb{C}^{M\times M}$, or equivalently, its scattering matrix $\mathbf{\Phi}\in\mathbb{C}^{M\times M}$ [25]. Specifically, the former characterizes the relationship between voltages and currents at the $M$ ports of the reconfigurable admittance network, while the latter characterizes the scattering behavior relating the incident and reflected waves. According to microwave network theory [4], we can map the scattering matrix $\bm{\Phi}$ to the admittance matrix $\mathbf{Y}$ by
@@
where $Y_{0}$ is the characteristic admittance and usually set as $Y_{0}=\frac{1}{50}$ siemens (S). Depending on the circuit topology of the $M$-port reconfigurable admittance network, the scattering matrix $\bm{\Phi}$ and admittance matrix $\mathbf{Y}$ exhibit different mathematical characteristics. In this paper, we consider two representative architectures of BD-RIS: the group-connected architecture [5] and the forest-connected architecture [10]. The detailed descriptions of these two architectures are shown below.
@@ II-A.1 Group-Connected Reconfigurable Admittance Network
In this architecture, the $M$ ports of the reconfigurable admittance network is uniformly divided into $G$ groups. The ports within the same group are all connected to each other via tunable admittance components, whereas the ports in different group are not connected to each other [5]. Fig. 1(a) shows a 36-element BD-RIS with the group-connected architecture of group size 3, where one group is randomly selected for illustration. For a group-connected reconfigurable admittance network with $G$ uniform groups, each group contains $\bar{M}=\frac{M}{G}$ elements, referred to as the group size. For the $g$-th group, all ports are connected with each other, leading to a full admittance matrix $\mathbf{Y}_{g}\in\mathbb{C}^{\bar{M}\times\bar{M}}$. Hence, the admittance matrix $\mathbf{Y}$ has a block-diagonal structure given by
@@
For a reciprocal reconfigurable admittance network, $\mathbf{Y}_{g},$ $\forall g\in\mathcal{G}=\{1,\ldots,G\}$ is symmetric, that is
$\mathbf{Y}_{g}=\mathbf{Y}^{T}_{g},\forall g\in\mathcal{G}.$ (3)
More specifically, as illustrated in Fig. 1(a), in the $g$-th group of the reconfigurable admittance network, each port $m_{g}=(g-1)\bar{M}+m$ is connected to port $n_{g}=(g-1)\bar{M}+n$ through a tunable admittance $Y_{m_{g},n_{g}},\forall m,n\in\bar{\mathcal{M}}=\{1,\ldots,\bar{M}\}$. In addition, each port $m_{g}$ is also connected to ground via an admittance $Y_{m_{g},m_{g}}$. According to [4], $\mathbf{Y}_{g}$ is written as
$[\mathbf{Y}_{g}]_{m,n}=\begin{cases}-Y_{m_{g},n_{g}},&m_{g}\neq n_{g};\\ \sum_{k=(g-1)\bar{M}+1}^{g\bar{M}}Y_{m_{g},k},&m_{g}=n_{g}.\end{cases}$ (4)
The circuit complexity of group-connected BD-RIS, defined as the required number of tunable admittance components, is given by $\frac{M(\bar{M}+1)}{2}$. An extreme case is the fully-connected architecture [5], with $\bar{M}=M$, and thus a circuit complexity of $\frac{M(M+1)}{2}$. The circuit complexity of the group-connected architecture can be significantly high, particularly when the group size $\bar{M}$ is large. To reduce this complexity while maintaining the flexibility of the reconfigurable admittance network, the forest-connected architecture for BD-RIS has been proposed in [10].

### v2

@@ Abstract
Beyond diagonal reconfigurable intelligent surface (BD-RIS) has emerged as an advancement and generalization of the conventional diagonal RIS (D-RIS) by introducing tunable interconnections between RIS elements, enabling smarter wave manipulation and enlarged coverage. While BD-RIS has demonstrated advantages over D-RIS in various aspects, most existing works rely on the assumption of a lossless model, leaving practical considerations unaddressed. This paper thus proposes a lossy BD-RIS model and develops corresponding optimization algorithms for various BD-RIS-aided communication systems. First, by leveraging admittance parameter analysis, we model each tunable admittance component based on a lumped circuit with losses and derive an expression of a circle characterizing the real and imaginary parts of each tunable admittance. We then consider the received signal power maximization in single-user single-input single-output (SISO) systems with the proposed lossy BD-RIS model. To solve the formulated challenging optimization problem, we design an effective algorithm by carefully exploiting the problem structure. In particular, an alternating direction method of multipliers (ADMM) framework is custom-designed to deal with the complicated constraints associated with lossy BD-RIS. Furthermore, we extend the proposed algorithmic framework to more general multiuser multiple-input single-output (MU-MISO) systems, where the transmit precoder and BD-RIS scattering matrix are jointly designed to maximize the sum-rate of the system. Finally, simulation results demonstrate that all BD-RIS architectures still outperform D-RIS in the presence of losses, but the optimal BD-RIS architectures in the lossless case are not necessarily optimal in the lossy case, e.g., group-connected BD-RIS can outperform fully- and tree-connected BD-RISs in SISO systems with relatively high losses at BD-RIS, whereas the opposite always holds true in the lossless case.
@@ I Introduction
Reconfigurable intelligent surface (RIS) is a nearly-passive planar surface composed of numerous reflecting elements, each capable of inducing a phase and/or amplitude change to the impinging wave. This flexibility enables dynamic wave manipulation, which provides enhanced channel gain and enlarges wireless coverage [smart_radio_Marco_Di_renzo, IRS_magazine_Qingqing_Wu, IRS_tutorial_Qingqing_Wu]. According to microwave network theory [pozar2009microwave], RIS can be generally modeled as numerous scattering elements connected to a multiport reconfigurable impedance network [shen2021modelling]. Most existing studies focus on RIS designs with a single-connected architecture, where each port of the reconfigurable impedance network is connected to the ground via a tunable impedance. This mathematically results in a diagonal scattering matrix [shen2021modelling], and such a single-connected RIS is also referred to as diagonal (D)-RIS. Recently, a novel branch of RIS, known as beyond diagonal (BD)-RIS [BD-RIS_magazine], has been proposed as a generalization of D-RIS. BD-RIS operates by introducing additional tunable impedance components that interconnect the ports in the reconfigurable impedance network, mathematically resulting in a scattering matrix that is no longer restricted to being diagonal. This enhanced flexibility, provided by those interconnections, enables a smarter wave manipulation, leading to greater channel gain, higher spectral efficiency, and broader wireless coverage compared to D-RIS [BD-RIS_magazine].
@@
Existing studies on BD-RIS have shown its advantages over D-RIS in various aspects. These works can be categorized as follows: [ios_hongliang_zhang, hybrid_hongyu, multisector_hongyu] focus on mode analysis; [shen2021modelling, nerinigraphtheory, dynamic_grouping_hongyu, band_stem_zheyu, non_reciprocal_hongyu] focus on architecture design; [close_form_nerini, channel_estimation_hongyu, universal_optimization_zheyu] focus on signal processing; [discrete_nerini, mutual_coupling_hongyu, mutual_coupling_nerini, lossy_interconnections_nerini, wideband_circuit_hongyu] focus on hardware impairments. However, most existing works [shen2021modelling, ios_hongliang_zhang, hybrid_hongyu, multisector_hongyu, nerinigraphtheory, dynamic_grouping_hongyu, close_form_nerini, channel_estimation_hongyu, non_reciprocal_hongyu, universal_optimization_zheyu, discrete_nerini, mutual_coupling_hongyu, mutual_coupling_nerini, wideband_circuit_hongyu, band_stem_zheyu] assume the reconfigurable impedance network is lossless to simplify the theoretical analysis and explore the performance upperbound of BD-RIS architectures. One exception work [lossy_interconnections_nerini] models BD-RIS with lossy interconnections in the reconfigurable impedance network using transmission line theory, where these interconnections are in series with tunable impedance components. However, the tunable impedance components themselves are still assumed to be lossless. In practice, semiconductor devices such as positive-intrinsic-negative (PIN) diodes and varactor diodes are used to implement the tunable impedance components. These semiconductor devices have inherent resistances, which dissipate power in the circuit. Therefore, assuming a reconfigurable impedance network with lossless tunable components results in a model that underestimates the performance degradation in realistic systems. It is crucial to analyze the characteristics of BD-RIS with lossy components in the reconfigurable impedance network and establish an accurate model that reflects these losses.
@@
In the works on practical D-RIS modeling [practicalRIS_Abeywickrama, practical_ris_wenhao, practical_mimo_ofdm_ris_hongyu], a lumped circuit model for each RIS element has been shown to effectively capture the relationship between reflection amplitude and phase. However, generalizing this approach to lossy BD-RIS is challenging. D-RISs are typically modeled using scattering parameters, where each entry in the scattering matrix represents the amplitude and phase response of an individual RIS element, allowing them to be modeled independently. In contrast, for BD-RISs, the losses of BD-RIS are embedded in each lossy reconfigurable component, and the interconnections between RIS elements cause the amplitude and phase of each entry in the scattering matrix to depend on other entries as well. In this sense, it is impossible to model lossy BD-RIS by independently characterizing its entries in the scattering matrix. Consequently, while lossless BD-RIS can be modeled directly by a beyond-diagonal scattering matrix, which is an upgrade of lossless D-RIS, the lossy BD-RIS needs an accurate model to reflect the losses embedded in the circuit, which, unfortunately, cannot be directly shown in the scattering matrix. In [universal_framework_nerini], a universal framework has been proposed that unifies the analysis of scattering, impedance, and admittance parameters for characterizing RIS-aided models, and presents the advantages of each type of parameter analysis in RIS modeling and optimization. While the aforementioned works [shen2021modelling, discrete_nerini, close_form_nerini, hybrid_hongyu, multisector_hongyu, nerinigraphtheory, dynamic_grouping_hongyu, channel_estimation_hongyu, non_reciprocal_hongyu] primarily focus on modeling and designing BD-RIS-aided systems using scattering parameter, admittance parameter has proven to be a powerful approach in characterizing the circuit topology models of BD-RIS [lossy_interconnections_nerini, wideband_circuit_hongyu, universal_optimization_zheyu]. Admittance parameter analysis enables the modeling of BD-RIS with lossy interconnections using transmission line theory [lossy_interconnections_nerini], the wideband modeling of BD-RIS through circuit-based approaches [wideband_circuit_hongyu], and the development of an efficient architecture-independent optimization framework for BD-RIS in general multiuser systems [universal_optimization_zheyu].
Inspired by the advantages of admittance parameter analysis as in [lossy_interconnections_nerini, wideband_circuit_hongyu, universal_optimization_zheyu] to model and optimize BD-RIS, we propose to model lossy BD-RIS by individually accounting for each lossy tunable component by its admittance parameter. The contributions of this work are summarized as follows.
@@
First, we propose a novel model for BD-RIS that incorporates losses in the reconfigurable admittance network. Specifically, by utilizing admittance parameter analysis, we represent the lossy reconfigurable admittance network by modeling each tunable admittance component individually using a lumped circuit model based on a practical varactor. In the proposed model, the real and imaginary parts of each admittance component form a circle, with feasible values constrained to a segment of the circle determined by the practical range of the tunable capacitance. The proposed model is sufficiently general to be applicable to any BD-RIS architectures, such as fully-/group-connected [shen2021modelling] and tree-/forest-connected architectures [^1] The proposed model can also be applied to band-connected and stem-connected architectures [band_stem_zheyu]. [nerinigraphtheory].
@@
Second, we optimize the scattering matrix of BD-RIS to maximize the received signal power for a BD-RIS-aided single-user single-input single-output (SISO) system, incorporating the proposed lossy BD-RIS model. To solve this problem, we employ a minorization-maximization ($M$) framework to transform the problem into a more tractable form. To deal with the complicated constraints introduced by lossy BD-RIS, we further custom-design an alternating direction method of multipliers (ADMM) algorithm to solve the resulting subproblems. In addition, to reduce the computational complexity, we propose a low-complexity solution that minimizes the Euclidean distance of an admittance matrix condition derived from the lossless BD-RIS model, using a similar ADMM algorithm to handle the constraints introduced by lossy BD-RIS.
@@
Fourth, we present simulation results to evaluate the performance of the proposed algorithms. Results show that all BD-RIS architectures outperform D-RIS under the lossy RIS modeling. In contrast to the lossless case where fully- and tree-connected BD-RISs always outperform group-connected BD-RIS in SISO systems, in the lossy case, group-connected BD-RIS can outperform fully- and tree-connected BD-RISs. Except for the fully-connected architecture, group-connected BD-RIS consistently outperforms forest-connected BD-RIS under the same group size [^2] Recall that in lossless BD-RIS, forest-connected and group-connected architectures achieve the same performance for a given group size in SISO systems [nerinigraphtheory, parteo_frontier_nerini], though the latter has higher complexity.. In MU-MISO systems, increasing interconnections does not always lead into better performance as well. Consequently, the Pareto frontier characterized for lossless BD-RIS [parteo_frontier_nerini, band_stem_zheyu] does not hold in the lossy case.
@@
Notations: Boldface lowercase and uppercase letters represent column vectors and matrices, respectively, while scalars are denoted with letters not in bold font. $\mathbb{R}$ and $\mathbb{C}$ indicate the set of real numbers and complex numbers, respectively. $(\cdot)^{*}$, $(\cdot)^{T}$, $(\cdot)^{H}$, and $(\cdot)^{-1}$ represent the conjugate, transpose, conjugate transpose, and inverse, respectively. $\Re\{\cdot\}$ and $\Im\{\cdot\}$ indicate the real and imaginary parts of a complex number, respectively. $\mathbb{E}\{\cdot\}$ denotes the statistical expectation. $|\cdot|$, $\|\cdot\|_{2}$ and $\|\cdot\|_{F}$ refer to the absolute-value norm, the $\ell_{2}$ norm, and the Frobenius norm, respectively. $\arg(\cdot)$ denotes the phase of a complex number. $\otimes$ represents the Kronecker product. $\mathsf{blkdiag}(\cdot)$ represents a block-diagonal matrix. $\mathsf{vec}(\cdot)$ denotes the vectorization of a matrix, $\overline{\mathsf{vec}}(\cdot)$ is the reverse operation of vectorization, and $\mathsf{tr}(\cdot)$ denotes the trace of a matrix. $\mathbf{I}_{M}$ denotes an $M\times M$ identity matrix. $\jmath=\sqrt{-1}$ represents the imaginary unit. $\mathbf{A}]_{i,j}$ and $\mathbf{a}]_{i}$ indicate the $(i,j)$-th entry of $\mathbf{A}$ and the $i$-th entry of $\mathbf{a}$, respectively. Finally, $\mathbf{A}]_{i:i^{\prime},j:j^{\prime}}$ represents the submatrix of $\mathbf{A}$ formed by selecting the $i$-th to $i^{\prime}$-th rows and $j$-th to $j^{\prime}$-th columns.
@@ II-A Admittance Parameter Analysis of BD-RIS
An $M$-element RIS is a passive device, which is modeled as $M$ elements connected to an $M$-port reconfigurable admittance network [shen2021modelling]. The reconfigurable admittance network, composed of tunable passive components, can be mathematically described by its admittance matrix $\mathbf{Y}\in\mathbb{C}^{M\times M}$, or equivalently, its scattering matrix $\mathbf{\Phi}\in\mathbb{C}^{M\times M}$ [universal_framework_nerini]. Specifically, the former characterizes the relationship between voltages and currents at the $M$ ports of the reconfigurable admittance network, while the latter characterizes the scattering behavior relating the incident and reflected waves. According to microwave network theory [pozar2009microwave], we can map the scattering matrix $\bm{\Phi}$ to the admittance matrix $\mathbf{Y}$ by
@@
where $Y_{0}$ is the characteristic admittance and usually set as $Y_{0}=\frac{1}{50}$ siemens (S). Depending on the circuit topology of the $M$-port reconfigurable admittance network, the scattering matrix $\bm{\Phi}$ and admittance matrix $\mathbf{Y}$ exhibit different mathematical characteristics. In this paper, we consider two representative architectures of BD-RIS: the group-connected architecture [shen2021modelling] and the forest-connected architecture [nerinigraphtheory]. The detailed descriptions of these two architectures are shown below.
@@ II-A.1 Group-Connected Reconfigurable Admittance Network
In this architecture, the $M$ ports of the reconfigurable admittance network is uniformly divided into $G$ groups. The ports within the same group are all connected to each other via tunable admittance components, whereas the ports in different group are not connected to each other [shen2021modelling]. Fig. 1(a) shows a 36-element BD-RIS with the group-connected architecture of group size 3, where one group is randomly selected for illustration. For a group-connected reconfigurable admittance network with $G$ uniform groups, each group contains $\bar{M}=\frac{M}{G}$ elements, referred to as the group size. For the $g$-th group, all ports are connected with each other, leading to a full admittance matrix $\mathbf{Y}_{g}\in\mathbb{C}^{\bar{M}\times\bar{M}}$. Hence, the admittance matrix $\mathbf{Y}$ has a block-diagonal structure given by
@@
For a reciprocal reconfigurable admittance network, $\mathbf{Y}_{g},~\forall g\in\mathcal{G}=\{1,\ldots,G\}$ is symmetric, that is
$\mathbf{Y}_{g}=\mathbf{Y}^{T}_{g},~\forall g\in\mathcal{G}.$ (3)
More specifically, as illustrated in Fig. 1(a), in the $g$-th group of the reconfigurable admittance network, each port $m_{g}=(g-1)\bar{M}+m$ is connected to port $n_{g}=(g-1)\bar{M}+n$ through a tunable admittance $Y_{m_{g},n_{g}},~\forall m,n\in\bar{\mathcal{M}}=\{1,\ldots,\bar{M}\}$. In addition, each port $m_{g}$ is also connected to ground via an admittance $Y_{m_{g},m_{g}}$. According to [pozar2009microwave], $\mathbf{Y}_{g}$ is written as
$\mathbf{Y}_{g}]_{m,n}=\begin{cases}-Y_{m_{g},n_{g}},&m_{g}\neq n_{g};\\ \sum_{k=(g-1)\bar{M}+1}^{g\bar{M}}Y_{m_{g},k},&m_{g}=n_{g}.\end{cases}$ (4)
The circuit complexity of group-connected BD-RIS, defined as the required number of tunable admittance components, is given by $\frac{M(\bar{M}+1)}{2}$. An extreme case is the fully-connected architecture [shen2021modelling], with $\bar{M}=M$, and thus a circuit complexity of $\frac{M(M+1)}{2}$. The circuit complexity of the group-connected architecture can be significantly high, particularly when the group size $\bar{M}$ is large. To reduce this complexity while maintaining the flexibility of the reconfigurable admittance network, the forest-connected architecture for BD-RIS has been proposed in [nerinigraphtheory].